OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. @Mathematics

OpenStudy (aroub):

parallel=same slope y=mx+b , m=-1 , (-4,2) 2=-1(-4)+b 2=4+b -2=b so, y=-x-2